# Provably Safe Reinforcement Learning using Entropy Regularizer

### v2

@@ 6 Conclusion and Future Works
$\begin{split}&\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\\ &\overset{\overset{}{<}}{\sim}\frac{|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}}{p-p^{s}}+\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}\\ &+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}+\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &\overset{\overset{(a)}{<}}{\sim}\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}\end{split}$ (35)
@@
$\begin{split}&\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\\ &\overset{\overset{}{<}}{\sim}\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\end{split}$ (36)
@@
$\begin{split}R(K)&\overset{<}{\sim}\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\\ &\overset{<}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}\right)\end{split}$
@@
$\tilde{V}^{\tilde{P}_{k}}_{\pi_{k}}(c,x_{0})\leq V^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})$
@@
$\begin{split}V^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})&\overset{}{=}V^{\tilde{P}_{k}}_{\pi_{k}}({c},x_{0})-\frac{4T_{max}}{p-p^{s}}\sum_{(x,a)\in\mathcal{X}\times\mathcal{A}}\tilde{\xi}^{*}(x,a)\hat{\epsilon}_{k}(x,a)\end{split}$ (38)
From (37) and (38), if $\alpha=\frac{8T^{2}_{max}}{\text{log}(\frac{2T_{max}}{T_{max}+\eta})(p-p^{s})}$, we get $\tilde{V}^{\tilde{P}_{k}}_{\pi_{k}}(c,x_{0})\leq{V}^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})$.
@@
We get inequality $(a)$ in view of (39). Inequality $(b)$ is from the definition of $\tilde{V}^{\tilde{P}_{k}^{2}}_{\pi_{k}^{2}}(c,x_{0})$. Relation $(c)$ follows from the fact that $H(\tilde{\xi}^{*}(x,a))\leq\text{log}\left(\frac{2T_{max}}{\eta}\right)\frac{\tilde{\xi}^{*}(x,a)}{2T_{max}}$. Relation $(d)$ is due to Lemma 23. To get the final relation, we use (27) in the proof of Theorem 11 and Lemma 21.

### v3

@@ 6 Conclusion and Future Works
$\begin{split}&\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\\ &\overset{\overset{}{<}}{\sim}\frac{|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}}{p-p^{s}}+\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}\\ &+|\mathcal{X}|^{2}|\mathcal{A}|T^{2}_{max}+\frac{T^{3}_{max}|\mathcal{X}|^{2}|\mathcal{A}|}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\\ &+\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &\overset{\overset{(a)}{<}}{\sim}\sqrt{\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}}\Big(\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\Big)^{\frac{1}{2}}\\ &+\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}.\end{split}$ (36)
@@
$\begin{split}&\sum_{k}\Big(V^{P^{\tau}}_{\pi_{k}}(c,x_{0})-V^{\tilde{P}_{k}^{\tau}}_{\pi_{k}}(\tilde{c}_{k},x_{0})\Big)\\ &\overset{\overset{}{<}}{\sim}\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}.\end{split}$ (37)
@@
$\begin{split}R(K)&\overset{<}{\sim}\frac{\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}}{p-p^{s}}+\frac{|\mathcal{X}|^{3}|\mathcal{A}|T^{4}_{max}L}{p-p^{s}}\\ &\overset{<}{\sim}\tilde{\mathcal{O}}\left(\frac{1}{p-p^{s}}\sqrt{|\mathcal{X}|^{2}|\mathcal{A}|T^{3}_{max}K}\right).\qed\end{split}$
@@
$\tilde{V}^{\tilde{P}_{k}}_{\pi_{k}}(c,x_{0})\leq V^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0}).$
@@
$\begin{split}V^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})&\overset{}{=}V^{\tilde{P}_{k}}_{\pi_{k}}({c},x_{0})-\frac{4T_{max}}{p-p^{s}}\sum_{(x,a)\in\mathcal{X}\times\mathcal{A}}\tilde{\xi}^{*}(x,a)\hat{\epsilon}_{k}(x,a).\end{split}$ (39)
From (38) and (39), if $\alpha=\frac{8T^{2}_{max}}{\text{log}(\frac{2T_{max}}{T_{max}+\eta})(p-p^{s})}$, we get $\tilde{V}^{\tilde{P}_{k}}_{\pi_{k}}(c,x_{0})\leq{V}^{\tilde{P}_{k}}_{\pi_{k}}(\tilde{c}_{k},x_{0})$. ∎
@@
We get inequality $(a)$ in view of (40). Inequality $(b)$ is from the definition of $\tilde{V}^{\tilde{P}_{k}^{2}}_{\pi_{k}^{2}}(c,x_{0})$. Relation $(c)$ follows from the fact that $H(\tilde{\xi}^{*}(x,a))\leq\text{log}\left(\frac{2T_{max}}{\eta}\right)\frac{\tilde{\xi}^{*}(x,a)}{2T_{max}}$. $\tilde{P}^{2,\tau}_{k}$ is defined according to (20). Relation $(d)$ is due to Lemma 23. To get the final relation, we use (28) in the proof of Theorem 11 and Lemma 21. ∎